# Geographic Variation in

### v1

@@ 6 Interpreting Regional Differences in Mobility
Why does intergenerational mobility vary across regions? Following Chetty et al. (2014), the literature has searched for regional correlates that can explain some of the regional variation in mobility. This literature has been fruitful, in that it has identified various factors associated with low mobility, which can then be studied further in more targeted, causal research designs.
@@
To provide a simple example, imagine that income mobility might vary across regions either because the returns to human capital differ (e.g., because of different industrial composition) or because the intergenerational transmission of human capital differs (e.g., because of differences in residential segregation and the quality of the schools and teachers; see e.g. Durlauf, 1996). In a simple theoretical model of transmission, the latter would have more persistent effects on the prospects of families with low human capital (as we show below). By comparing regional differences in inter- and multigenerational correlations researchers might be able to distinguish these two stories.
@@ 6.1 A simple latent factor model
illustrating that any further decay of status advantages over more than two generations depends only on the “transferability” of endowments $\lambda_{r}$ and not on their returns $\rho_{r}$. In other words, different components of the transmission process affect inter- and multigenerational correlations differently, and their comparison therefore tells us something new about the source of mobility differences across regions. [^33] Our argument here is an example for the more general insight that the observation of the dynamics of an outcome variable can be helpful in parameter identification of its underlying model (Aigner et al., 1984).
@@
as is the case for most Swedish regions (see Section 3). As this finding of “excess persistence” $\Delta>0$ is common, the model has been used in several recent multigenerational studies (e.g., Braun and Stuhler, 2018; Neidhöfer and Stockhausen, 2019; Belloc et al., 2024; Celhay and Gallegos, 2025).
@@
Of course, this model is highly stylized, abstracting from many mechanisms that one might want to study explicitly. Its one-parent structure abstracts from the role of assortative mating, although the parameter $\lambda_{r}$ can be interpreted as a reduced-form representation of the combined effects of intergenerational transmission and assortative mating (Braun and Stuhler, 2018). It also ignores that the intergenerational transmission process may differ by gender (e.g., Pembrey et al., 2006) or neighborhood or peer effects. Perhaps most importantly, we assumed that the underlying distributional moments are in a steady state, which is implausible for data spanning three generations (especially given the large shifts in means shown in Table 1). In our application, we abstract from shifts in the variance across generations by focusing on correlation rather than regression coefficients. Moreover, we define $\beta_{-1}$ as an appropriately weighted average of the correlations between children and parents (G1-G2) and between parents and grandparents (G2-G3). [^34] Specifically, $\beta_{-1}=\sqrt{\beta_{-1}^{G1-G2}\beta_{-1}^{G2-G3}}$. See Atkinson and Jenkins (1984) and Nybom and Stuhler (2019) for a discussion of steady-state assumptions in mobility research. Braun and Stuhler (2018) discuss the estimation of a latent factor model with time-varying parameters.
@@
Still, the model captures the notion that an individual’s observed outcome $y$ might be a poor proxy for other endowments and advantages that influence the prospects of one’s child; and it shows that this simple notion has important implications for the relative size of inter- and multigenerational correlations. Extended versions of this model can explain the pattern of intergenerational transmission across different kins (Collado, Ortuño-Ortín and Stuhler, 2023) or rationalize why intergenerational transmission is higher at the surname than the family level (Clark, 2014).
@@
One obvious extension would be to allow for direct effects of parental outcomes $y_{t-1}$ on child outcomes $y_{t}$ in eq. (2), as in Becker and Tomes (1986) or Solon (2004). However, a commonly used version of this model as described in Solon (2004) implies that multigenerational correlations should be smaller than one might expect from parent-child correlations, i.e. $\Delta<0$, and is therefore less consistent with the observed multigenerational patterns. [^35] Put differently, this model implies that in a regression of child on parent and grandparent status, the coefficient on grandparents is negative; see Solon (2014) for details. And while studying more general versions of the Becker-Tomes model would be interesting, we do not pursue this path here, as with three generations there is not sufficient empirical information to identify the parameters of such extended models (see Collado, Ortuño-Ortín and Stuhler, 2023, for a more general model fitted on more extensive set of kinship moments).
@@ 6.4 Robustness and limitations
We next test the robustness of our results to different sample restrictions and empirical specifications. Table C.3 in the Appendix replicates our main analysis (Table 5) in a balanced sample restricted to complete child-parent-grandparent lineages. The results remain similar. In Table C.4, we show that they also remain similar if rather than years of schooling we use an indicator for high/low schooling (defined relative to the median schooling of the respective generation). This suggests that our main conclusions are not sensitive to the specific way how educational outcomes are measured, or to the fact that educational outcomes vary less in the earliest generation. Our conclusions are also robust to considering the log-level rather than the level of the correlations and parameters. [^38] As from equation (5) it follows that $log(\beta_{-2,r})=2log(\rho_{r})+2log(\lambda_{r})$ we may prefer regressing the log of the grandfather-child correlation in years of schooling on the log of the estimated returns $\rho_{r}$ and transferability $\lambda_{r}$, to compare their relative contribution to the regression $R^{2}$. We do so in Appendix Table C.5, which confirms our results from Table 5: variation in $\beta_{-2,r}$ is primarily due to the variation in $\lambda_{r}$ rather than variation in $\rho_{r}$. In fact, a simpler and more direct way of seeing this is by comparing the variance of $\hat{\lambda}$ and $\hat{\rho}$ across regions; for education (income), the variance of $log(\hat{\lambda})$ is nearly six (nearly four) times larger than the variance of $log(\hat{\rho})$ And while the inter- and multigenerational correlations differ by gender, we find a similar pattern when repeating the analysis using only male or only female children (Appendix Table C.2).
@@ 7 Conclusions
Intergenerational mobility varies substantially across regions, and this variability may be informative about the mechanisms by which advantages are being transmitted from parents to children. In this paper, we studied how the strength of multigenerational correlations varies across municipalities in Sweden, and argued that these multigenerational patterns provide additional insights on why parent-child correlations differ across regions. Our analysis therefore aimed to combine the advantages of two recent strands of the literature on geographic (e.g., Chetty et al., 2014) and multigenerational patterns in mobility (e.g., Lindahl et al., 2015). We also linked to a recent debate on the relation between mobility and cross-sectional inequality (e.g., Durlauf, Kourtellos and Tan, 2022).
@@
As a starting point, we first described the extent of regional variation in educational and earnings mobility across Swedish municipalities, and its spatial pattern. Educational mobility tends to be lower in the more densely populated southern parts of Sweden, and in and around the major cities. The pattern of earnings mobility is more dispersed, but upward mobility tends to be comparatively high in many municipalities in Southern Sweden, for children born around Stockholm, and in the northernmost municipalities in Sweden. We also studied whether the regional rankings are stable to the choice of mobility measures. We find that different measures of educational mobility are highly correlated, and similar for father-child as for mother-child correlations, and for inter- (parent-child) vs. multigenerational (grandparent-child) correlations. But regional rankings are more sensitive to the choice of measure for earnings mobility, and areas that are characterized by high mobility in earnings are not necessarily the same areas that are characterized by high mobility in education.
These regional variations are interesting from a descriptive perspective, but our main purpose was to exploit them to learn more about the intergenerational transmission process, and the relation between cross-sectional inequality and intergenerational mobility. We find that municipalities characterized by high earnings inequality in the parent generation also tend to be characterized by low intergenerational mobility. This relationship holds for most mobility measures that we considered, but the magnitudes vary considerably. One interesting observation is that the inequality-mobility relation is particularly pronounced for intergenerational correlations in education, in particular multigenerational correlations.
@@
Finally, we demonstrated that a comparison of inter- and multigenerational correlations can provide insights into how and why mobility varies across regions. Interpreting those regional variations through the lens of a latent factor model, we showed that the regional differences in mobility appear to be due to differences in the transmission process from parents to children itself, rather than variation in how latent endowments translate into observed advantages. A similar argument can be made with regard to the determinants of cross-sectional inequality. While our findings illustrate the potential value of combining regional and multigenerational perspectives on intergenerational mobility, a more detailed examination would likely offer additional insights.
@@ Appendix A Supplementary Evidence for Sections 2 and 3
Notes: Density plot of the estimated intergenerational coefficients in earnings ranks for father-child pairs across 163 “small” municipalities ($\leq$ 2,000 observations), 127 “large” municipalities ($>$ 2,000 observations), and similar numbers of small and large “placebo” municipalities (random allocation of observations across municipalities).
@@
$\hat{\beta}_{f}^{edu}$ $\hat{\beta}_{fgf}^{edu}$ $\hat{\beta}_{m}^{edu}$ $\hat{\beta}_{mgf}^{edu}$ $\hat{\beta}_{f}^{rank}$ $\hat{\beta}_{fgf}^{rank}$ $\hat{\beta}_{m}^{rank}$ $\hat{\beta}_{mgf}^{rank}$ $\hat{\beta}_{f}^{IGE}$ $\hat{\beta}_{fgf}^{IGE}$ $\hat{\beta}_{m}^{IGE}$ $\hat{\beta}_{mgf}^{IGE}$ $\hat{\beta}_{f}^{edu}$ 1.00 $\hat{\beta}_{fgf}^{edu}$ 0.79 1.00 $\hat{\beta}_{m}^{edu}$ 0.83 0.69 1.00 $\hat{\beta}_{mgf}^{edu}$ 0.75 0.80 0.72 1.00 $\hat{\beta}_{f}^{rank}$ 0.06 -0.00 0.08 -0.01 1.00 $\hat{\beta}_{fgf}^{rank}$ 0.35 0.33 0.28 0.31 0.51 1.00 $\hat{\beta}_{m}^{rank}$ 0.05 0.04 0.18 0.09 0.57 0.35 1.00 $\hat{\beta}_{mgf}^{rank}$ 0.29 0.25 0.30 0.28 0.51 0.56 0.44 1.00 $\hat{\beta}_{f}^{IGE}$0.11 -0.21 -0.08 -0.20 0.91 0.39 0.49 0.38 1.00 $\hat{\beta}_{fgf}^{IGE}$ 0.23 0.19 0.17 0.17 0.54 0.94 0.39 0.54 0.47 1.00 $\hat{\beta}_{m}^{IGE}$0.03 -0.07 0.10 -0.02 0.55 0.29 0.95 0.36 0.55 0.35 1.00 $\hat{\beta}_{mgf}^{IGE}$ 0.14 0.08 0.16 0.12 0.55 0.53 0.48 0.94 0.47 0.56 0.43 1.00
Notes: Correlations between different estimates of municipality-specific inter- and multigenerational measures. We exclude municipalities with fewer than 1000 observations and weight the correlation by sample size. Subscript $f$ stands for fathers (e.g., $\hat{\beta}_{f}^{edu}$ is the father-child correlation in years of schooling), $m$ for mothers, $fgf$ for paternal grandfathers, $mgf$ for maternal grandfathers.
@@
Dependent variable: Grandfather-child correlation $\beta_{-2,r}$ Panel A: Sons Years of schooling Earnings ranks (1) (2) (3) (4) (5) (6) $\hat{\rho}_{r}$ (“returns”) -0.176*** 0.086*** -$0$001 (0.023) (0.023) (0.004) (0.000) $\hat{\lambda}_{r}$ (“transferability”) 0.335*** 0.476*** 0.193*** 0.195*** (0.016) (0.027) (0.009) (0.012) # Municipalities 275 290 275 272 289 272 Adj. $R^{2}$ 0.386 0.849 0.901 0.084 0.811 0.762 Panel B: Daughters Dependent variable: Grandfather-child correlation $\beta_{-2,r}$ Years of schooling Earnings ranks (1) (2) (3) (4) (5) (6) $\hat{\rho}_{r}$ (“returns”) -0.130*** 0.049** -0.080*** 0.008 (0.013) (0.017) (0.016) (0.010) $\hat{\lambda}_{r}$ (“transferability”) 0.270*** 0.369*** 0.167*** 0.173*** (0.012) (0.033) (0.009) (0.015) # Municipalities 268 290 268 288 289 287 Adj. $R^{2}$ 0.399 0.823 0.821 0.272 0.665 0.656
@@
Notes: The dependent variable is the log grandfather-child correlation in years of schooling (columns 1-3) or the log grandfather-child correlation in earnings ranks (columns 4-6) across 290 Swedish municipalities. The independent variables are $log(\hat{\rho})$ and $log(\hat{\lambda})$, see equations (6) and (7). Birth cohorts 1981-89, weighted by the number of grandfather-child pairs in each municipality. Robust standard errors in parentheses:* p<0.05, ** p<0.01, *** p<0.005.

### v2

@@ 6 Interpreting Regional Differences in Mobility
Why does intergenerational mobility vary across regions? Following CHKS_Where_QJE, the literature has searched for regional correlates that can explain some of the regional variation in mobility. This literature has been fruitful, in that it has identified various factors associated with low mobility, which can then be studied further in more targeted, causal research designs.
@@
To provide a simple example, imagine that income mobility might vary across regions either because the returns to human capital differ (e.g., because of different industrial composition) or because the intergenerational transmission of human capital differs (e.g., because of differences in residential segregation and the quality of the schools and teachers; see e.g. Durlauf1996). In a simple theoretical model of transmission, the latter would have more persistent effects on the prospects of families with low human capital (as we show below). By comparing regional differences in inter- and multigenerational correlations researchers might be able to distinguish these two stories.
@@ 6.1 A simple latent factor model
illustrating that any further decay of status advantages over more than two generations depends only on the “transferability” of endowments $\lambda_{r}$ and not on their returns $\rho_{r}$. In other words, different components of the transmission process affect inter- and multigenerational correlations differently, and their comparison therefore tells us something new about the source of mobility differences across regions. [^33] Our argument here is an example for the more general insight that the observation of the dynamics of an outcome variable can be helpful in parameter identification of its underlying model (AignerHsiaoKapteynWansbeek1984).
@@
as is the case for most Swedish regions (see Section 3). As this finding of “excess persistence” $\Delta>0$ is common, the model has been used in several recent multigenerational studies (e.g., ECOJ:ECOJ12453; NeidhoeferStockhausen; Bellocetal2024; Celhaygallegos2025).
@@
Of course, this model is highly stylized, abstracting from many mechanisms that one might want to study explicitly. Its one-parent structure abstracts from the role of assortative mating, although the parameter $\lambda_{r}$ can be interpreted as a reduced-form representation of the combined effects of intergenerational transmission and assortative mating (ECOJ:ECOJ12453). It also ignores that the intergenerational transmission process may differ by gender (e.g., pembrey2006sex) or neighborhood or peer effects. Perhaps most importantly, we assumed that the underlying distributional moments are in a steady state, which is implausible for data spanning three generations (especially given the large shifts in means shown in Table 1). In our application, we abstract from shifts in the variance across generations by focusing on correlation rather than regression coefficients. Moreover, we define $\beta_{-1}$ as an appropriately weighted average of the correlations between children and parents (G1-G2) and between parents and grandparents (G2-G3). [^34] Specifically, $\beta_{-1}=\sqrt{\beta_{-1}^{G1-G2}\beta_{-1}^{G2-G3}}$. See AtkinsonJenkins1984 and NybomStuhler2019 for a discussion of steady-state assumptions in mobility research. ECOJ:ECOJ12453 discuss the estimation of a latent factor model with time-varying parameters.
@@
Still, the model captures the notion that an individual’s observed outcome $y$ might be a poor proxy for other endowments and advantages that influence the prospects of one’s child; and it shows that this simple notion has important implications for the relative size of inter- and multigenerational correlations. Extended versions of this model can explain the pattern of intergenerational transmission across different kins (ColladoOrtunoStuhlerKinship) or rationalize why intergenerational transmission is higher at the surname than the family level (Clark2014book).
@@
One obvious extension would be to allow for direct effects of parental outcomes $y_{t-1}$ on child outcomes $y_{t}$ in eq. (2), as in becker1986human or solon2004model. However, a commonly used version of this model as described in solon2004model implies that multigenerational correlations should be smaller than one might expect from parent-child correlations, i.e. $\Delta<0$, and is therefore less consistent with the observed multigenerational patterns. [^35] Put differently, this model implies that in a regression of child on parent and grandparent status, the coefficient on grandparents is negative; see Solon201413 for details. And while studying more general versions of the Becker-Tomes model would be interesting, we do not pursue this path here, as with three generations there is not sufficient empirical information to identify the parameters of such extended models (see ColladoOrtunoStuhlerKinship, for a more general model fitted on more extensive set of kinship moments).
@@ 6.4 Robustness and limitations
[…]
@@ 7 Conclusions
Intergenerational mobility varies substantially across regions, and this variability may be informative about the mechanisms by which advantages are passed from parents to children. In this paper, we studied how multigenerational correlations vary across municipalities in Sweden, and argued that these multigenerational patterns provide additional insights on why parent-child correlations differ across regions. Our analysis therefore aimed to combine the advantages of two recent strands of the literature on geographic (e.g., CHKS_Where_QJE) and multigenerational patterns in mobility (e.g., LindahlPalme2014_IGE4Generations). We also linked to a recent debate on the relation between mobility and cross-sectional inequality (e.g., durlauf2022great).
@@
As a starting point, we described the extent of regional variations in inter- and multigenerational educational and earnings mobility across Swedish municipalities, and their spatial patterns. Educational mobility tends to be lower in the more densely populated southern parts of Sweden and around major cities, while the pattern of earnings mobility is more dispersed. These regional rankings are largely stable to the choice of mobility measures. Different measures of educational mobility are highly correlated, and similar for inter- (parent-child) and multigenerational (grandparent-child) correlations along both the paternal and maternal lines. Regions also rank similarly across inter- and multigenerational correlations in earnings ranks, although these patterns are more dispersed. Moreover, areas with high earnings mobility are not necessarily those with high educational mobility.
These regional variations are interesting from a descriptive perspective, but our main purpose was to exploit them to learn more about intergenerational processes. We found that regional grandfather-child correlations are consistently larger than the square of their parent-child counterparts, implying that the latter only capture part of the intergenerational process, and tend to understate status transmission in the “long run” (across many generations). But while understating the level of transmission, parent-child correlations still capture differences in long-run transmission across regions. Conventional mobility statistics thus remain useful for comparative purposes, irrespectively of whether our interest centers on two- or multigenerational processes.
We further replicate the well-established pattern that regions characterized by high earnings inequality also tend to be characterized by low mobility, including across three generations. This relationship holds for most mobility measures that we considered, but the magnitudes vary. One interesting observation is that the inequality-mobility relation is particularly pronounced for educational mobility, especially for multigenerational correlations in education.
@@
Finally, we illustrated that a comparison of inter- and multigenerational correlations can provide insights into how and why mobility differs across regions. Interpreting those regional differences through the lens of a latent factor model, we showed that they appear to be due to variation in the transmission of latent advantages from parents to children, rather than from differences in how those latent advantages convert into observable outcomes. A similar argument can be made with regard to the determinants of cross-sectional inequality. While these findings are stylized, they illustrate the potential value of integrating regional and multigenerational perspectives on intergenerational mobility.
@@ Appendix A Supplementary Evidence for Sections 2 and 3
Notes: Density plot of the estimated intergenerational coefficients in earnings ranks for father-child pairs across 163 “small” municipalities ($\leq$ 2,000 observations), 127 “large” municipalities (>> 2,000 observations), and similar numbers of small and large “placebo” municipalities (random allocation of observations across municipalities).
@@
$\hat{\beta}_{f}^{edu}$ $\hat{\beta}_{fgf}^{edu}$ $\hat{\beta}_{m}^{edu}$ $\hat{\beta}_{mgf}^{edu}$ $\hat{\beta}_{f}^{rank}$ $\hat{\beta}_{fgf}^{rank}$ $\hat{\beta}_{m}^{rank}$ $\hat{\beta}_{mgf}^{rank}$ $\hat{\beta}_{f}^{IGE}$ $\hat{\beta}_{fgf}^{IGE}$ $\hat{\beta}_{m}^{IGE}$ $\hat{\beta}_{mgf}^{IGE}$ $\hat{\beta}_{f}^{edu}$ 1.00 $\hat{\beta}_{fgf}^{edu}$ 0.79 1.00 $\hat{\beta}_{m}^{edu}$ 0.83 0.69 1.00 $\hat{\beta}_{mgf}^{edu}$ 0.75 0.80 0.72 1.00 $\hat{\beta}_{f}^{rank}$ 0.06 -0.00 0.08 -0.01 1.00 $\hat{\beta}_{fgf}^{rank}$ 0.35 0.33 0.28 0.31 0.51 1.00 $\hat{\beta}_{m}^{rank}$ 0.05 0.04 0.18 0.09 0.57 0.35 1.00 $\hat{\beta}_{mgf}^{rank}$ 0.29 0.25 0.30 0.28 0.51 0.56 0.44 1.00 $\hat{\beta}_{f}^{IGE}$ -0.11 -0.21 -0.08 -0.20 0.91 0.39 0.49 0.38 1.00 $\hat{\beta}_{fgf}^{IGE}$ 0.23 0.19 0.17 0.17 0.54 0.94 0.39 0.54 0.47 1.00 $\hat{\beta}_{m}^{IGE}$ -0.03 -0.07 0.10 -0.02 0.55 0.29 0.95 0.36 0.55 0.35 1.00 $\hat{\beta}_{mgf}^{IGE}$ 0.14 0.08 0.16 0.12 0.55 0.53 0.48 0.94 0.47 0.56 0.43 1.00
Notes: Correlations between different estimates of municipality-specific inter- and multigenerational measures. We exclude municipalities with fewer than 1000 observations and weight the correlation by sample size. Subscript $f$ stands for fathers (e.g., $\hat{\beta}_{f}^{edu}$ is the father-child correlation in years of schooling), $m$ for mothers, fgffgf for paternal grandfathers, mgfmgf for maternal grandfathers.
@@
Dependent variable: Grandfather-child correlation $\beta_{-2,r}$ Panel A: Sons Years of schooling Earnings ranks (1) (2) (3) (4) (5) (6) $\hat{\rho}_{r}$ (“returns”) -0.176*** 0.086*** -0.010* -0.001 (0.023) (0.023) (0.004) (0.000) $\hat{\lambda}_{r}$ (“transferability”) 0.335*** 0.476*** 0.193*** 0.195*** (0.016) (0.027) (0.009) (0.012) # Municipalities 275 290 275 272 289 272 Adj. $R^{2}$ 0.386 0.849 0.901 0.084 0.811 0.762 Panel B: Daughters Dependent variable: Grandfather-child correlation $\beta_{-2,r}$ Years of schooling Earnings ranks (1) (2) (3) (4) (5) (6) $\hat{\rho}_{r}$ (“returns”) -0.130*** 0.049** -0.080*** 0.008 (0.013) (0.017) (0.016) (0.010) $\hat{\lambda}_{r}$ (“transferability”) 0.270*** 0.369*** 0.167*** 0.173*** (0.012) (0.033) (0.009) (0.015) # Municipalities 268 290 268 288 289 287 Adj. $R^{2}$ 0.399 0.823 0.821 0.272 0.665 0.656
@@
[…]